# Satisficing Exploration in Bandit Optimization

### v1

@@ 4 Regret Analysis
Under Assumption 1, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
@@
We also show that, in the non-realizable case, SELECT enjoys the same regret bound as the oracle in Assumption 1 (The proof is provided in Section B of the appendix).
@@
Under Assumption 1, if $r(X^{*})<S$, the standard regret of SELECT is bounded by
@@
With the above results, we establish that under a very general assumption, SELECT does achieve constant satisficing regret in the realizable case without compromising the standard regret guarantee in the non-realizable case. Altogether, these empower SELECT the potential to become a general algorithm for decision-making under uncertainty.
@@ 5 Examples
In what follows, we instantiate SELECT to several popular bandit optimization settings, including finite-armed bandits, concave bandits, and Lipschitz bandits. Along the way, we showcase how our results enable constant satisficing regret for bandits with large and even infinite arm space, which was not achieved by existing algorithms.
1. Finite-Armed Bandits:
Consider a $K$-armed bandit problem, i.e., $\mathcal{X}=[K]$. In this case, both the UCB algorithm (see, e.g., Bubeck and Cesa-Bianchi (2012)) and Thompson sampling (see, e.g., Agrawal and Goyal (2017)) achieve a regret bound of $O(\sqrt{KT\log(T)})$. Combining them with Theorems 1 and 2, we have the following corollary. At the end of this section, we also provide some discussions on the lower bounds of the satisficing regrets.
@@ 1. Finite-Armed Bandits:
By using either the UCB algorithm or Thompson sampling as $\texttt{ALG}(\cdot)$, if $r(X^{*})\geq S$, the satisficing regret of SELECT is bounded by
@@
If $r(X^{*})<S$, then the standard regret of SELECT is bounded by
$\texttt{Regret}\leq\sqrt{KT}\cdot\textup{polylog}(T).$
Remark 3 (Comparison with Garivier et al. (2019), Michel et al. (2023)).
Garivier et al. (2019), Michel et al. (2023) also provide algorithms for satisficing regret for finite-armed bandits. While our satisficing regret bound is incomparable to Garivier et al. (2019), which attains $O(K/\Delta_{S})$ satisficing regret, we provide major improvement over Michel et al. (2023), which achieves $O(K/\Delta_{S}+K/(\Delta_{S}^{*})^{2})$ satisficing regret. Compared to Michel et al. (2023), our regret bound remove the dependence on $\Delta_{S}$ and the additional $1/\Delta_{S}^{*}$ factor. We also point out that if we want to achieve the same satisficing regret as Garivier et al. (2019), we can simply change the LCB test in Step 3 to a UCB test.
2. Concave Bandits:
Consider a concave bandit setting, i.e., $\mathcal{X}\subset\mathcal{R}^{d}$ is a convex set, and $r(X)$ is a concave and $1$-Lipschitz continuous function defined on $\mathcal{X}$. Agarwal et al. (2011) gives an algorithm that enjoys $\text{poly}(d)\sqrt{T}\cdot\text{polylog}(T)$ regret. Together with Theorems 1 and 2, we have the following corollary.
@@ 2. Concave Bandits:
By using the algorithm in Agarwal et al. (2011) as $\texttt{ALG}(\cdot)$, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
@@
If $r(X^{*})<S$, then the standard regret of SELECT is bounded by
$\texttt{Regret}\leq\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T).$
3. Lipschitz Bandits:
Consider a Lipschitz bandit setting in $d$ dimensions, i.e., $\mathcal{X}=[0,1]^{d}$, and $r(X)$ is an $L$-Lipschitz function. Slivkins et al. (2019) (see Section 4.2) introduces a UCB algorithm to achieve a $O(L^{1/(d+2)}T^{(d+1)/(d+2)}\log(T)^{1/(d+2)})$ regret upper bound bound. Together with Theorem 1 and Theorem 2, we have the following corollary.
@@ 3. Lipschitz Bandits:
By using the UCB algorithm with uniform discretization in Slivkins et al. (2019) as $\texttt{ALG}(\cdot)$, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{\dfrac{L}{(\Delta_{S}^{*}/2)^{d+1}}\cdot\textup{polylog}\left(\dfrac{L}{\Delta_{S}^{*}}\right),\,L^{\frac{1}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T)\right\};$
If $r(X^{*})<S$, then the standard regret of SELECT is bounded by
$\texttt{Regret}\leq L^{\frac{1}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T).$
@@
4. Lower Bounds:
To complement our main results, we also present the satisficing regret lower bounds for finite-armed bandits and concave bandits. The first one is a satisficing regret lower bound for the finite-armed bandits.
@@ 6.1 Finite-Armed Bandits
Setup: In this case, we consider an instance of $4$ arms, and the expected rewards of all arms are $\{0.6,0.7,0.8,1\}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We conduct experiments for both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.93$; for the non-realizable case, we set the satisficing level $S=1.5$. In both cases, we compare SELECT against Thompson sampling in Agrawal and Goyal (2017) as well as SAT-UCB and SAT-UCB+ in Hajiabolhassan and Ortner (2023). The experiment is repeated for 1000 times and we report the average results.
Results: The results of the realizable case is presented in Figure 2(a), and the results of the non-realizable case is presented in Figure 2(b). From Figure 2(a) one can see that SELECT, SAT-UCB and SAT-UCB+ exhibit constant satisficing regret, and incur less satisficing regret compared to Thompson sampling. Furthermore, SELECT incurs less regret compared to SAT-UCB and achieves comparable performance with SAT-UCB+ in the realizable case. From Figure 2(b) one can see that SELECT incurs roughly the same standard regret as Thompson sampling. Furthermore, SELECT also incurs less standard regret in the non-realizable case compared to either SAT-UCB or SAT-UCB+.
@@ 6.2 Concave Bandits
Setup: In this case, we consider an instance with arm set $[0,1]$ and concave reward function $r(x)=1-16(x-0.25)^{2}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.3$; for the non-realizable case, we set the satisficing level $S=1.5$. For both cases, we compare our algorithm SELECT with the convex bandit algorithm introduced in Agarwal et al. (2011). We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm space uniformly with stepsize $L^{1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm space. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case is provided in Figure 3(a), and the results of non-realizable case is provided in Figure 3(b). From Figure 3(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than the algorithm in Agarwal et al. (2011), SAT-UCB and SAT-UCB+. The reason why SAT-UCB and SAT-UCB+ has not converged is that they usually repeatedly pull non-satisficing arms which are close to satisficing level but provide empirically highest UCB. As shown in Figure 3(b), SELECT can also reach good enough distribution compared to the algorithm in Agarwal et al. (2011) , SAT-UCB, and SAT-UCB+ in the non-realizable case.
@@ 6.3 Lipschitz Bandits
Setup: In this case, we consider a two-dimensional Lipschitz bandit in domain $(x,y)\in[0,1]^{2}$ with Lipschitz rewards $f(x,y)=\min\{1,3e^{-100((x-0.5)^{2}+(y-0.7)^{2})}\}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level S = 0.5; for the non-realizable case, we set the satisficing level S = 1.5. For both cases, we compare our algorithm SELECT with the uniformly discretized UCB introduced in Slivkins et al. (2019). We again use SAT-UCB and SAT-UCB+ as heuristics by uniformly discretizing the arm space with stepsize $L^{1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and run the algorithms over the discretized arms. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case is provided in Figure 4(a), and the results of non-realizable case is provided in Figure 4(b). From Figure 4(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than uniform UCB, SAT-UCB and SAT-UCB+. From 4(b), we can see that SELECT still works the best while Uniform UCB, SAT-UCB and SAT-UCB+ have similar performance.
@@ 7 Conclusion
In this paper, we propose SELECT, a general algorithmic template, for satisficing exploration in bandit optimization. For a given bandit optimization problem and a sub-linear regret learning oracle, we show that SELECT attains constant satisficing regret in the realizable case and the same regret as the learning oracle in the non-realizable case. We instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and also make a discussion on the corresponding lower bounds. Finally, numerical experiments show that SELECT attains constant regret in several popular bandit problems and demonstrates excellent performances when compared to other benchmarks in both the realizable and the non-realizable cases.
@@ Appendix A Proof of Theorem 1
$\dfrac{1}{\delta_{0}}=\dfrac{C_{1}K_{\alpha,\beta}\log(1/\delta_{0})^{\beta^{\prime}}}{\Delta_{S}^{*}}=\dfrac{(4\beta^{\prime})^{\beta^{\prime}}C_{1}C_{\alpha,\beta}\log(2/\delta_{0}^{1/4\beta^{\prime}})^{\beta^{\prime}}}{\Delta_{S}^{*}}\leq\dfrac{(8\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}}{\Delta_{S}^{*}(\delta_{0})^{1/4}},$
@@ A.2 Proof of Proposition 2
By Assumption 1 and recall from the proof of Proposition 1, we have
@@
According to the assumption of this proposition, we have $i$ satisfies the condition

### v2

@@ 4 Regret Analysis
If $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
@@
We also show that, in the non-realizable case, SELECT enjoys the same regret bound as the oracle in Condition 1 (The proof is provided in Section B of the appendix).
@@
If $r(X^{*})<S$, the standard regret of SELECT is bounded by
@@
With the above results, we establish that under a very general condition, SELECT does achieve constant satisficing regret in the realizable case without compromising the standard regret guarantee in the non-realizable case. Altogether, these empower SELECT the potential to become a general algorithm for decision-making under uncertainty.
@@ 5 Examples
In what follows, we instantiate SELECT to several popular bandit optimization settings, including finite-armed bandits, concave bandits, and Lipschitz bandits. Along the way, we showcase how our results enable constant satisficing regret for bandits with large and even infinite arm space, which was not achieved by existing algorithms. We remark that the examples given here are non-exhaustive and similar results can be derived for other bandit optimization settings such as linear bandits. In all three examples of this section, we assume the mean reward of each arm is in $[0,1]$.
1. Finite-Armed Bandits: Consider a finite-armed bandit problem with $K$ arms, i.e., $\mathcal{X}=[K]$. In this case, both the UCB algorithm (see, e.g., Bubeck and Cesa-Bianchi 2012) and Thompson sampling (see, e.g., Agrawal and Goyal 2017) achieve a regret bound of $O(\sqrt{KT\log(T)})$. Combining them with Theorems 1 and 2, we have the following corollary. At the end of this section, we also provide some discussions on the lower bounds of the satisficing regrets.
@@
By using either the UCB algorithm or Thompson sampling as ALG, if $r(X^{*})\geq S$, the satisficing regret of SELECT is bounded by
@@
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq\sqrt{KT}\cdot\textup{polylog}(T).$
Remark 3 (Comparison with Garivier et al. 2019, Michel et al. 2023).
Garivier et al. (2019), Michel et al. (2023) also provide algorithms for satisficing regret for finite-armed bandits. While our satisficing regret bound is incomparable to Garivier et al. (2019), which attains a $O(K/\Delta_{S})$ satisficing regret, we provide major improvement over Michel et al. (2023), which achieves $O(K/\Delta_{S}+K/(\Delta_{S}^{*})^{2})$ satisficing regret. Compared to Michel et al. (2023), our regret bound remove the dependence on $\Delta_{S}$ and the additional $1/\Delta_{S}^{*}$ factor. We also point out that if we want to achieve the same satisficing regret as Garivier et al. (2019), we can simply change the LCB test in Step 3 to a UCB test.
2. Concave Bandits: Consider a concave bandit setting, i.e., $\mathcal{X}\subset\mathbb{R}^{d}$ is a bounded convex set, and $r(X)$ is a concave and $1$-Lipschitz continuous function defined on $\mathcal{X}$. Agarwal et al. (2011) gives an algorithm that enjoys $\text{poly}(d)\sqrt{T}\cdot\text{polylog}(T)$ regret. Together with Theorems 1 and 2, we have the following corollary.
@@
By using the algorithm in Agarwal et al. (2011) as ALG, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
@@
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T).$
3. Lipschitz Bandits: Consider a Lipschitz bandit setting in $d$ dimensions, i.e., $\mathcal{X}=[0,1]^{d}$, and $r(X)$ is an $L$-Lipschitz function (here Lipschitz functions are defined in the sense of $\infty$-norm). Bubeck et al. (2011) introduces a uniformly discretized UCB algorithm to achieve a $O(L^{d/(d+2)}T^{(d+1)/(d+2)}\sqrt{\log(T)})$ regret upper bound bound. Together with Theorem 1 and Theorem 2, we have the following corollary.
@@
By using the UCB algorithm with uniform discretization in Bubeck et al. (2011) as ALG, if $r(X^{*})\geq S$, then the satisficing regret of SELECT is bounded by
$\texttt{Regret}_{S}\leq\min\left\{\dfrac{L^{d}}{(\Delta_{S}^{*}/2)^{d+1}}\cdot\textup{polylog}\left(\dfrac{L}{\Delta_{S}^{*}}\right),\,L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T)\right\};$
If $r(X^{*})<S$, the standard regret of SELECT is bounded by $\texttt{Regret}\leq L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T).$
@@
4. Lower Bounds: To complement our main results, we also present the satisficing regret lower bounds for finite-armed bandits and concave bandits. The first one is a satisficing regret lower bound for the finite-armed bandits.
@@ 6.1 Finite-Armed Bandits
Setup: In this case, we consider an instance of $4$ arms, and the expected rewards of all arms are $\{0.6,0.7,0.8,1\}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We conduct experiments for both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.93$; for the non-realizable case, we set the satisficing level $S=1.5$. In both cases, we compare SELECT against Thompson sampling in Agrawal and Goyal (2017) as well as SAT-UCB and SAT-UCB+ in Michel et al. (2023). For SELECT, we use Thompson Sampling as the learning oracle. The experiment is repeated for 1000 times and we report the average results.
Results: The results of the realizable case are presented in Figure 2(a), and the results of the non-realizable case are presented in Figure 2(b). From Figure 2(a) one can see that SELECT, SAT-UCB and SAT-UCB+ exhibit constant satisficing regret, and incur less satisficing regret compared to Thompson sampling. Furthermore, SELECT incurs less regret compared to SAT-UCB and achieves comparable performance with SAT-UCB+ in the realizable case. From Figure 2(b) one can see that SELECT incurs roughly the same standard regret as Thompson sampling. Furthermore, SELECT also incurs less standard regret in the non-realizable case compared to either SAT-UCB or SAT-UCB+.
@@ 6.2 Concave Bandits
Setup: In this case, we consider an instance with arm set $[0,1]$ and concave reward function $r(x)=1-16(x-0.25)^{2}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.3$; for the non-realizable case, we set the satisficing level $S=1.5$. For both cases, we compare our algorithm SELECT with the convex bandit algorithm introduced in Agarwal et al. (2011). For SELECT, we use the algorithm in Agarwal et al. (2011) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm space uniformly with stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm space. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 3(a), and the results of non-realizable case are provided in Figure 3(b). From Figure 3(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than the algorithm in Agarwal et al. (2011), SAT-UCB and SAT-UCB+. The reason why SAT-UCB and SAT-UCB+ has not converged is that they usually repeatedly pull non-satisficing arms which are close to satisficing level but provide empirically highest UCB. As shown in Figure 3(b), SELECT can also reach good enough distribution compared to the algorithm in Agarwal et al. (2011) , SAT-UCB, and SAT-UCB+ in the non-realizable case.
@@ 6.3 Lipschitz Bandits
Setup: In this case, we consider a two-dimensional Lipschitz bandit in domain $(x,y)\in[0,1]^{2}$ with Lipschitz rewards $f(x,y)=\min\{1,3e^{-100((x-0.5)^{2}+(y-0.7)^{2})}\}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level S = 0.5; for the non-realizable case, we set the satisficing level S = 1.5. For both cases, we compare our algorithm SELECT with the uniformly discretized UCB introduced in Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”). For SELECT, we use the Uniform UCB as the learning oracle. We again use SAT-UCB and SAT-UCB+ as heuristics by uniformly discretizing the arm space with stepsize $L^{-1/4}T^{-1/4}$, where $L$ is the Lipschitz constant of the reward function, and run the algorithms over the discretized arms. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 4(a), and the results of non-realizable case are provided in Figure 4(b). From Figure 4(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than Uniform UCB, SAT-UCB and SAT-UCB+. From 4(b), we can see that SELECT still works the best while Uniform UCB, SAT-UCB and SAT-UCB+ have similar performance.
@@ 7 Conclusion
In this paper, we propose SELECT, a general algorithmic template, for satisficing regret minimization in bandit optimization. For a given bandit optimization problem and a sub-linear regret learning oracle, we show that SELECT attains constant satisficing regret in the realizable case and the same regret as the learning oracle in the non-realizable case. We instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and also make a discussion on the corresponding lower bounds. Finally, numerical experiments show that SELECT attains constant regret in several popular bandit problems and demonstrates excellent performances when compared to other benchmarks in both the realizable and the non-realizable cases.
For future studies, we suggest further exploration in the following directions. First, our work mainly focuses on the stationary settings, and further research could explore satisficing in the nonstationary online learning settings (Wei and Luo 2021, Cheung et al. 2022). Also, the notion of satisficing regret indeed discourages exploration. Therefore, another direction is to study the role of satisficing regret in settings with limited adaptivity (Gao et al. 2019, Feng et al. 2024).
Appendix
@@ Appendix A Proof of Theorem 1
$\dfrac{1}{\delta_{0}}=\dfrac{C_{1}K_{\alpha,\beta}\log(1/\delta_{0})^{\beta^{\prime}}}{\Delta_{S}^{*}}=\dfrac{(4\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}\log(2/\delta_{0}^{1/4\beta^{\prime}})^{\beta^{\prime}}}{\Delta_{S}^{*}}\leq\dfrac{(8\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}}{\Delta_{S}^{*}(\delta_{0})^{1/4}},$
@@ A.1 Proof of Proposition 1
$\left\lceil\dfrac{1}{\gamma_{i}^{2}}\right\rceil\cdot\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-1}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
@@ A.2 Proof of Proposition 2
By Condition 1 and recall from the proof of Proposition 1, we have
@@
According to the condition of this proposition, we have $i$ satisfies the condition
@@ C.2 Proof of Theorem 4
Appendix D Choices of $C_{1},\alpha,\beta$ for Several Bandit Problem Classes
In this section, we provide a detailed discussion on the choice of $C_{1},\alpha,\beta$ for several bandit problem classes. In three cases
D.1 Finite-armed Bandits
Consider any instance of finite-armed bandits with mean reward bounded by $[0,1]$. By Theorem 7.2 of Lattimore and Szepesvári (2020), for any given time horizon $t\geq 2$, the standard regret of the UCB algorithm is upper bounded by
$8\sqrt{Kt\log(t)}+3\sum_{k=1}^{K}(r(k^{*})-r(k))\leq 8\sqrt{Kt\log(t)}+3K,$
where $K\geq 2$ is the number of arms and $k^{*}$ is the optimal arm.
When $t>K$, the regret is upper bounded by $8\sqrt{Kt\log(t)}+3K\leq 11\sqrt{Kt\log(t)}$, where the last inequality is due to $t>K\geq 2$, thus $t\geq 3$ and $\log(t)\geq\log(3)>1$; When $t\leq K$, the regret is upper bounded by $t\leq\sqrt{Kt}\leq 11\sqrt{Kt\log(2)}\leq 11\sqrt{Kt\log(t)}$.
Thus for any given time horizon $t\geq 2$, the regret of the UCB algorithm is bounded by $11\sqrt{Kt\log(t)}$. Therefore, by setting $C_{1}=11\sqrt{K}$ and $\alpha=\beta=1/2$ and using UCB algorithm as the oracle, Condition 1 is satisfied for finite-armed bandits.
D.2 Concave Bandits
Consider any concave bandits with $\mathcal{X}\subseteq\mathbb{R}^{d}$ being a compact and convex set contained in a Euclidean ball with radius $1$, and reward function being concave and $1$-Lipschitz. By Theorem 2 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 2 in Agarwal et al. (2011) is upper bounded by
$768d^{3}\sqrt{t}\log(t)^{2}\left(\dfrac{2d^{2}\log(d)}{c_{2}^{2}}+1\right)\left(\dfrac{4d^{7}c_{1}}{c_{2}^{3}}+\dfrac{d(d+1)}{c_{2}}\right)\left(\dfrac{4c_{1}d^{4}}{c_{2}^{2}}+11\right),$
where $c_{1},c_{2}$ are algorithm hyperparameters that satisfies $c_{1}\geq 64$, $c_{2}\leq 1/32$. Therefore Condition 1 is satisfied for concave bandits by using Algorithm 2 in Agarwal et al. (2011) and setting $\alpha=1/2$, $\beta=2$ and
$C_{1}=768d^{3}\left(\dfrac{2d^{2}\log(d)}{c_{2}^{2}}+1\right)\left(\dfrac{4d^{7}c_{1}}{c_{2}^{3}}+\dfrac{d(d+1)}{c_{2}}\right)\left(\dfrac{4c_{1}d^{4}}{c_{2}^{2}}+11\right),$
Furthermore, if the instance of concave bandit is one-dimensional, i.e., $\mathcal{X}=[0,1]$, by Theorem 1 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 1 in Agarwal et al. (2011) is bounded by
$108\sqrt{t\log(t)}\cdot\log_{4/3}(t)=108/\log(4/3)\sqrt{t}\cdot\log(t)^{3/2}.$
Therefore Condition 1 is also satisfied for one-dimensional concave bandits by using Algorithm 1 in Agarwal et al. (2011) and setting $C_{1}=108/\log(4/3)$, $\alpha=1/2$, $\beta=3/2$.
D.3 Lipschitz Bandits
Consider any Lipschitz bandits with arm set being $[0,1]^{d}$, reward function being $L$-Lipschitz and mean reward bounded on $[0,1]$ (here we define Lipschitz continuity in the sense of $\infty$-norm). By Section 2.2 of Bubeck et al. (2011), if the regret of the UCB algorithm for finite-armed bandits with $K$ arms is bounded by $c_{\text{ALG}}\sqrt{Kt\log(t)}$ for any $t\geq 2$, then for any given time horizon $t\geq 2$, the standard regret of the uniformly discretized UCB algorithm introduced by Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”) is upper bounded by
$(1+c_{\text{ALG}})L^{d/(d+2)}t^{(d+1)/(d+2)}\cdot\sqrt{\log(t)}.$
Using the results in Section D.1, we have $c_{\text{ALG}}=11$, thus the regret of the Uniform UCB algorithm is bounded by
$12L^{d/(d+2)}t^{(d+1)/(d+2)}\cdot\sqrt{\log(t)}.$
Therefore, by setting $C_{1}=12L^{d/(d+2)}$, $\alpha=(d+1)/(d+2)$, $\beta=1/2$ and using Uniform UCB as the oracle, Condition 1 is satisfied for Lipschitz bandits.
Appendix E Ablation Study for SELECT
In this section, we use an ablation study to demonstrate that all three steps of SELECT plays a vital role in obtaining a constant regret bound. Specifically, we compare the performance of SELECT with the following adapted versions of SELECT:
  • SELECT without step 1: instead of identifying candidate satisficing arms using a learning oracle, the candidate satisficing arm is drawn uniformly at random from the arm set before we proceed to forced sampling and LCB tests in each round;
  • SELECT without step 2: After identifying a candidate satisficing arm from step 1 in each round, the forced sampling in step 2 is skipped and a LCB test on the candidate satisficing arm is immediately started;
  • SELECT without step 3: Each round is terminated after running step 1 and step 2 without entering the LCB test in step 3.
Setup: We conduct the ablation test using an instance of the Lipschitz bandit. We use the reward function defined in Section 6.3, and set the satisficing threshold $S=0.7$. The learning oracle we use in step 1 is Uniform UCB introduced in Bubeck et al. (2011).
Results: The results are provided in Figure 5. One can see that after removing any of the three steps, our algorithm is unable to maintain constant satisficing regret in the realizable case. Specifically, step 1 serves the purpose of efficiently identifying a candidate satisficing arm. If step 1 is replaced by randomly drawing candidate satisficing arms from the arm set, then the satisficing regret is almost linear in $T$, indicating that the algorithm struggles to identify satisficing arms without the help of the learning oracle. Step 2 and step 3 combined mainly serve the purpose of exploiting the candidate satisficing arm and at the same time, testing whether it is truly satisficing. Without either step 2 or step 3, the satisficing regret starts a round of rapid increase whenever the time horizon $T$ exceeds certain thresholds, thus fail to maintain constant satisficing regret. This indicates that without either step 2 or step 3, the algorithm almost always terminates a round shortly after completing step 1 even if the candidate satisficing arm obtained from step 1 is indeed satisficing. We refer to Remark 2 for a detailed discussion on the reason for this. Therefore we conclude that all three steps of SELECT play a vital role in obtaining a constant satisficing regret bound in the realizable case.
Figure 5: Results of the ablation test
Appendix F Robustness on the choice of $\gamma_{i}$
(a) Realizable case
(b) Non-realizable case
Figure 6: Performance of SELECT with different choice of $\gamma_{i}$
In Algorithm 1, we define $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$ for all $i\in\mathbb{N}_{+}$, but all our results still hold if all $\gamma_{i}$ are multiplied by a constant. In this section, we conduct numerical experiments on testing the robustness of SELECT in the choice $\gamma_{i}$.
Setup: We conduct the experiment on the same instance of Lipschitz bandit in Section 6.3. We multiply the $\{\gamma_{i}\}$ in Algorithm 1 with a constant $\lambda$, and vary $\lambda=0.5,1,2,5$.
Results: The result of the realizable case is provided in Figure 6(a), and the result of the non-realizable case is provided in Figure 6(b). One can see from the results that different choices of $\gamma_{i}$ have limited impact on the empirical performance of SELECT, both in the realizable and non-realizable case. Therefore the empirical performance of SELECT is robust in the choice of $\gamma_{i}$.